# Gibbs-preserving operations requiring infinite amount of quantum coherence

### v2

@@ Appendix B Lower bound for pairwise reversible Gibbs-preserving Operations
where we used Theorem S.1 in the third line and (34) and (39) in the fourth line.
@@
Appendix C Construction of pairwise reversible Gibbs-preserving Operations
(Proof of Theorem 3)
@@ Appendix C Construction of pairwise reversible Gibbs-preserving Operations
$\langle i|\psi|i\rangle\neq 0,\quad\langle j|\psi|j\rangle\neq 0,\quad E_{i,S}\neq E_{j,S}$ (43)
@@
Let $\psi$ and $\phi$ be the pure states in (43) and (44). Define
$\Lambda(\rho)=\mathrm{Tr}(\psi\rho)\phi+\mathrm{Tr}\left[(\mathds{1}-\psi)\rho\right]\eta$ (45)
@@
$\eta=\frac{\tau_{S^{\prime}}-\mathrm{Tr}(\psi\tau_{S})\phi}{\mathrm{Tr}\left[(\mathds{1}-\psi)\tau_{S}\right]}.$ (46)
@@
$\phi\leq 2^{D_{\max}(\phi\|\tau_{S^{\prime}})}\tau_{S^{\prime}}=2^{D_{\min}(\psi\|\tau_{S})}\tau_{S^{\prime}}=\mathrm{Tr}(\psi\tau_{S})^{-1}\tau_{S^{\prime}}$ (47)
where the first inequality is by definition of $D_{\max}$, the first equality is due to (44), and the last equality is by definition of $D_{\min}$. This ensures that $\eta\geq 0$ and that $\Lambda$ is a valid Gibbs-preserving channel.
@@
$\mathrm{Tr}(\phi\eta)\propto\mathrm{Tr}(\phi\tau_{S^{\prime}})-\mathrm{Tr}(\psi\tau_{S})=2^{-D_{\min}(\phi\|\tau_{S^{\prime}})}-2^{-D_{\min}(\psi\|\tau_{S})}=0$ (48)
where the final equality is due to (44). This implies that $\phi$ and $\eta$ are perfectly distinguishable and thus ensures that $\delta(\Lambda,\mathbb{P})=0$ for a choice of state pair =$\mathbb{P}=\{\psi,\sigma\}$ where $\sigma$ is an arbitrary state in $S$ such that $\mathrm{Tr}(\psi\sigma)=0$.
As a choice of $\sigma$, we can particularly choose an orthogonal pure state $\psi^{\perp}$ defined as follows. Let $a_{i}\coloneqq\langle{i}|{\psi}\rangle$ and $a_{j}\coloneqq\langle{j}|{\psi}\rangle$ be the coefficients for $i$ th and $j$ th energy of $\psi$, where $i$ and $j$ are the labels in (43). We assume $a_{i,j}\in\mathbb{R}$ without loss of generality, as incoherent unitary—which can arbitrarily adjust the relative phase—is energy conserving. We then choose
$|\psi^{\perp}\rangle=\frac{1}{\sqrt{|a_{j}|^{2}+|a_{i}|^{2}}}\left(a_{j}|i\rangle-a_{i}|j\rangle\right),$ (49)
for which one can directly check that the condition $\mathrm{Tr}(\psi\psi^{\perp})=0$ is satisfied. Since
$\Lambda^{\dagger}(H_{S^{\prime}})=\mathrm{Tr}(\phi H_{S^{\prime}})\psi+\mathrm{Tr}(\eta H_{S^{\prime}})(\mathds{1}-\psi),$ (50)
@@
$\langle{\psi}|{\Lambda^{\dagger}(H_{S^{\prime}})}|{\psi^{\perp}}\rangle=0.$ (51)
@@
$\left|\langle{\psi}|{H-\Lambda^{\dagger}(H_{S^{\prime}})}|{\psi^{\perp}}\rangle\right|$ $=\left|\langle{\psi}|{H}|{\psi^{\perp}}\rangle\right|=\frac{|a_{i}a_{j}|}{\sqrt{a_{i}^{2}+a_{j}^{2}}}|E_{i,S}-E_{j,s}|>0$ (52)
@@
${\cal C}(\Lambda,\mathbb{P})=\left|\langle{\psi}|{H-\Lambda^{\dagger}(H_{S^{\prime}})}|{\psi^{\perp}}\rangle\right|>0$ (53)
@@
$\tau_{S^{\prime},i^{\prime}}=r\tau_{S,i}+(1-r)\tau_{S,j},$ (55)
whose existence is guaranteed because of (54). Let $\psi$ and $\phi$ be the pure states defined by
$|\psi\rangle$ $:=\sqrt{r}|i\rangle_{S}+\sqrt{1-r}|j\rangle_{S},$ (56)
$|\phi\rangle$ $:=|i^{\prime}\rangle_{S^{\prime}}.$ (57)
It suffices to show that $\psi$ and $\phi$ satisfy (43) and (44). It is straightforward to check (43) noting $10<r<$ 1. Eq. (44) can be checked as follows:
@@
$\Lambda(\rho)=\mathrm{Tr}(|{+}\rangle\!\langle{+}|_{i,i+1}\rho)|{0}\rangle\!\langle{0}|+\mathrm{Tr}(|{-}\rangle\!\langle{-}|_{i,i+1}\rho)|{1}\rangle\!\langle{1}|+\mathrm{Tr}(P_{i,i+1}^{\perp}\rho)\eta$ (59)
@@
$\eta\coloneqq\frac{\tau_{S^{\prime}}-\mathrm{Tr}(|{+}\rangle\!\langle{+}|_{i,i+1}\tau_{S})|{0}\rangle\!\langle{0}|-\mathrm{Tr}(|{-}\rangle\!\langle{-}|_{i,i+1}\tau_{S})|{1}\rangle\!\langle{1}|}{\mathrm{Tr}(P_{i,i+1}^{\perp}\tau_{S})}.$ (60)
@@
$\eta\propto\left(\tau_{S^{\prime},0}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\right)|{0}\rangle\!\langle{0}|+\left(\tau_{S^{\prime},1}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\right)|{1}\rangle\!\langle{1}|+\sum_{j\geq 2}\tau_{S^{\prime},j}|{j}\rangle\!\langle{j}|\geq 0$ (61)
@@
$\tau_{S^{\prime},k}-\tau_{S,k}=\frac{(Z-Z^{\prime})e^{-\beta E_{k}}}{ZZ^{\prime}}\geq 0,\quad Z\coloneqq\sum_{n=0}^{d-1}e^{-\beta E_{n}},Z^{\prime}\coloneqq\sum_{n=0}^{d^{\prime}-1}e^{-\beta E_{n}}$ (62)
@@
$\tau_{S^{\prime},0}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\geq\tau_{S,0}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\geq 0$ (63)
@@
$\Lambda^{\dagger}(H_{d^{\prime}})=\mathrm{Tr}(H_{d^{\prime}}|{0}\rangle\!\langle{0}|)|{+}\rangle\!\langle{+}|_{i,i+1}+\mathrm{Tr}(H_{d^{\prime}}|{1}\rangle\!\langle{1}|)|{-}\rangle\!\langle{-}|_{i,i+1}+\mathrm{Tr}(H_{d^{\prime}}\eta)P_{i,i+1}^{\perp}$ (64)
@@
+|${}_{i,i+1}\langle{+}|{\Lambda^{\dagger}(H_{d^{\prime}})}|{-}\rangle_{i,i+1}=0.$ (65)
@@
|+$\left|{}_{i,i+1}\langle{+}|{H_{d}}|{-}\rangle_{i,i+1}\right|=\frac{E_{i+1}-E_{i}}{2}>0$ (66)

### v3

@@ II.1 Proof of Theorem 1
where we used Theorem S.1 in the third line and (S.37) and (S.42) in the fourth line.
II.2 Extension to the cases without perfect pairwise reversibility
We here remark that we can extend Theorem S.2 to the case where $\Lambda$ is not pairwise reversible for $\mathbb{P}$.
Theorem S.3.
Let $\Lambda:S\to S^{\prime}$ be a Gibbs-preserving Operation. Then, for an arbitrary orthogonal pair states $\mathbb{P}$, the following relation holds:
$\sqrt{{\cal F}^{\epsilon}_{c}(\Lambda)}\geq\frac{{\cal C}(\Lambda,\mathbb{P})}{\epsilon+\delta(\Lambda,\mathbb{P})}-\Delta(H_{S})-3\Delta(H_{S^{\prime}}),$ (S.44)
The inequality (S.44) is a generalization of (S.35). Indeed, when $\Lambda$ is reversible for $\mathbb{P}$, (S.44) reduces to (S.35).
Proof.
Similarly to the proof of Theorem S.2, let $\Lambda_{\epsilon}$ be a channel that approximates $\Lambda$ with error $\epsilon$, i.e., $D_{F}(\Lambda_{\epsilon},\Lambda)\leq\epsilon$. We only have to show
$\delta(\Lambda_{\epsilon},\mathbb{P})\leq\delta(\Lambda,\mathbb{P})+\epsilon.$ (S.45)
The other parts are the same as the proof of Theorem S.2.
Let ${\cal R}$ be a recovery channel satisfying the following relation, whose existence is ensured by the definition of $\delta(\Lambda,\mathbb{P})$.
$\delta(\Lambda,\mathbb{P})=\frac{D_{F}({\cal R}\circ\Lambda(\psi),\psi)^{2}+D_{F}({\cal R}\circ\Lambda(\phi),\phi)^{2}}{2}.$ (S.46)
Then, we obtain (S.45) as follows:
$\delta(\Lambda_{\epsilon},\mathbb{P})^{2}$ $\leq\frac{D_{F}({\cal R}\circ\Lambda(\psi),\psi)^{2}+D_{F}({\cal R}\circ\Lambda(\phi),\phi)^{2}}{2}$
$\leq\frac{(D_{F}({\cal R}\circ\Lambda_{\epsilon}(\psi),\psi)+D_{F}({\cal R}\circ\Lambda_{\epsilon}(\psi),{\cal R}\circ\Lambda(\psi))^{2}+(D_{F}({\cal R}\circ\Lambda_{\epsilon}(\phi),\phi)+D_{F}({\cal R}\circ\Lambda_{\epsilon}(\phi),{\cal R}\circ\Lambda(\phi))^{2}}{2}$
$\leq\frac{(D_{F}({\cal R}\circ\Lambda_{\epsilon}(\psi),\psi)+\epsilon)^{2}+(D_{F}({\cal R}\circ\Lambda_{\epsilon}(\phi),\phi)+\epsilon)^{2}}{2}$
$=\frac{D_{F}({\cal R}\circ\Lambda_{\epsilon}(\psi),\psi)^{2}+2D_{F}({\cal R}\circ\Lambda_{\epsilon}(\psi),\psi)\epsilon+\epsilon^{2}+D_{F}({\cal R}\circ\Lambda_{\epsilon}(\phi),\phi)^{2}+2D_{F}({\cal R}\circ\Lambda_{\epsilon}(\phi),\phi)\epsilon+\epsilon^{2}}{2}$
$=\delta(\Lambda,\mathbb{P})^{2}+2\frac{D_{F}({\cal R}\circ\Lambda_{\epsilon}(\psi),\psi)+D_{F}({\cal R}\circ\Lambda_{\epsilon}(\phi),\phi)}{2}\epsilon+\epsilon^{2}$
$\leq\delta(\Lambda,\mathbb{P})^{2}+2\sqrt{\frac{D_{F}({\cal R}\circ\Lambda_{\epsilon}(\psi),\psi)^{2}+D_{F}({\cal R}\circ\Lambda_{\epsilon}(\phi),\phi)^{2}}{2}}\epsilon+\epsilon^{2}$
$=(\delta(\Lambda,\mathbb{P})+\epsilon)^{2}.$ (S.47)
II.3 Towards the extension to systems with unbounded Hamiltonians
So far, the results in this letter assume that $S$ and $S^{\prime}$ are finite dimensional systems, where $H_{S}$ and $H_{S^{\prime}}$ are guaranteed to be bounded operators. This assumption is crucial for Theorem 1 to be nontrivial, where the lower bound contains the quantities $\Delta(H_{S})$ and $\Delta(H_{S^{\prime}})$ that diverge for unbounded Hamiltonians. (We remark that Theorem 2 directly holds in this case as well because the proof does not rely on the spectra of $H_{S}$ and $H_{S^{\prime}}$.)
In this section, we provide a preliminary result toward extending Theorem 1 to the one that provides a nontrivial bound for $H_{S}$ and $H_{S^{\prime}}$ with unbounded spectra. The following result shows that the quantities $\Delta(H_{S})$ and $\Delta(H_{S^{\prime}})$ can be replaced by the ones that can take finite values even at the limit of unbounded Hamiltonians.
Theorem S.4.
For finite-dimensional systems $S$ and $S^{\prime}$, let $\Lambda:S\to S^{\prime}$ be a pairwise reversible Gibbs-preserving Operation with a reversible pure state pair :=$\mathbb{P}:=\{|{\psi}\rangle\!\langle{\psi}|,|{\phi}\rangle\!\langle{\phi}|\}$. Let $\tilde{\Lambda}:S\to S^{\prime}$ be a CPTP map satisfying $D_{F}(\Lambda,\tilde{\Lambda})\leq\epsilon$ and the following inequalities for a real positive number $\Delta$:
$\max_{\rho\in\mathrm{span}(\mathbb{P})}\left|\mathrm{Tr}\left(H_{S^{\prime}}\tilde{\Lambda}(\rho)\right)-\mathrm{Tr}\left(H_{S^{\prime}}\Lambda(\rho)\right)\right|$ $\leq\epsilon\Delta,$ (S.48)
$\max_{\rho\in\mathrm{span}(\mathbb{P})}|V(\tilde{\Lambda}(\rho);H_{S^{\prime}})-V(\Lambda(\rho);H_{S^{\prime}})|$ $\leq\epsilon\Delta^{2},$ (S.49)
where $\mathrm{span}(\mathbb{P})$ refers to the set of quantum states in the linear span of $\mathbb{P}$, and $V(\eta;O)$ denotes the variance of the observable $O$ for the state $\eta$. Then, the following inequality holds:
$\sqrt{{\cal F}_{c}(\tilde{\Lambda})}\geq\frac{{\cal C}(\Lambda,\mathbb{P})}{\epsilon}-2\max_{\rho\in\mathrm{span}(\mathbb{P})}(\sqrt{V(\rho;H_{S})}+\sqrt{V(\Lambda(\rho);H_{S^{\prime}})})-2(1+\sqrt{\epsilon})\Delta.$ (S.50)
Although this result assumes the finite dimensionality of $S$ and $S^{\prime}$, the inequality (S.50) does not directly depend on the maximum range of the spectra of $H_{S}$ or $H_{S^{\prime}}$. Therefore, it is expected that if we can extend the results in Ref. Tajima et al. (2022)—which assumes the finite dimensionality—to the infinite-dimensional cases involving unbounded operators, Theorem S.4 could also be extended to unbounded $H_{S}$ and $H_{S^{\prime}}$, which provides nontrivial lower bounds for coherence cost and diverges at the limit of $\epsilon\to 0$. We leave the thorough investigation of this problem as a future work.
Proof.
To prove (S.50), we use the following theorem:
Theorem S.5 (Ref. Tajima et al. (2022)).
For finite-dimensional systems $S$ and $S^{\prime}$, let ${\cal E}:S\to S^{\prime}$ and $\mathbb{P}$ be an arbitrary quantum channel and an arbitrary orthogonal state pair. Let $(U,\rho_{E},H_{E},H_{E^{\prime}})$ be an implementation of $\Lambda$ satisfying
$U^{\dagger}(H_{S^{\prime}}+H_{E^{\prime}})U$ $=H_{S}+H_{E}$ (S.51)
${\cal E}(...)$ $=\mathrm{Tr}_{E^{\prime}}[U...\otimes\rho_{E}U^{\dagger}].$ (S.52)
Then, it holds that
$\frac{{\cal C}({\cal E},\mathbb{P})}{\sqrt{{\cal F}(\rho_{E};H_{E})}+\Delta_{F}}\leq\delta({\cal E},\mathbb{P}),$ (S.53)
where
$\Delta_{F}:=\max_{\rho\in\mathrm{span}(\mathbb{P})}\sqrt{{\cal F}(\rho\otimes\rho_{E};H_{S}\otimes 1_{E}-U^{\dagger}H_{S^{\prime}}\otimes 1_{E^{\prime}}U)}$ (S.54)
with ${\cal F}(\eta;O)$ being the Fisher information of the observable $O$ for the state $\eta$.
One can upper bound $\Delta_{F}$ in Theorem S.5 by
$\Delta_{F}$ $=\max_{\rho\in\mathrm{span}(\mathbb{P})}\sqrt{{\cal F}(\rho\otimes\rho_{E};H_{S}\otimes 1_{E}-U^{\dagger}H_{S^{\prime}}\otimes 1_{E^{\prime}}U)}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\max_{\rho\in\mathrm{span}(\mathbb{P})}\left(\sqrt{{\cal F}(\rho\otimes\rho_{E};H_{S}\otimes 1_{E})}+\sqrt{{\cal F}(\rho\otimes\rho_{E};U^{\dagger}H_{S^{\prime}}\otimes 1_{E^{\prime}}U)}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2\max_{\rho\in\mathrm{span}(\mathbb{P})}\left(\sqrt{V(\rho\otimes\rho_{E};H_{S}\otimes 1_{E})}+\sqrt{V(\rho\otimes\rho_{E};U^{\dagger}H_{S^{\prime}}\otimes 1_{E^{\prime}}U)}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}2\max_{\rho\in\mathrm{span}(\mathbb{P})}\left(\sqrt{V(\rho;H_{S})}+\sqrt{V({\cal E}(\rho);H_{S^{\prime}})}\right),$ (S.55)
where we used $\sqrt{{\cal F}(\eta;O_{1}+O_{2})}\leq\sqrt{{\cal F}(\eta;O_{1})}+\sqrt{{\cal F}(\eta,O_{2})}$ in (a), ${\cal F}(\eta;O)\leq 4V(\eta;O)$ in (b) and $V(\rho\otimes\rho_{E};U^{\dagger}H_{S^{\prime}}\otimes 1_{E^{\prime}}U)=\mathrm{Tr}\left[U^{\dagger}H^{2}_{S^{\prime}}\otimes 1_{E^{\prime}}U(\rho\otimes\rho_{E})\right]-\mathrm{Tr}\left[U^{\dagger}H_{S^{\prime}}\otimes 1_{E^{\prime}}U(\rho\otimes\rho_{E})\right]^{2}=V({\cal E}(\rho);H_{S^{\prime}})$ in (c).
Also, when $(U,\rho_{E},H_{E},H_{E^{\prime}})$ implements a channel ${\cal E}$, we obtain
${\cal F}_{c}({\cal E})\geq{\cal F}_{c,{\rm cov}}({\cal E})\geq{\cal F}(\rho_{E};H_{E}).$ (S.56)
Combining (S.53), (S.55) and (S.56), and substituting $\tilde{\Lambda}$ for ${\cal E}$, we obtain
$\sqrt{{\cal F}_{c}(\tilde{\Lambda})}\geq\frac{{\cal C}(\tilde{\Lambda},\mathbb{P})}{\delta(\tilde{\Lambda},\mathbb{P})}-2\max_{\rho\in\mathrm{span}(\mathbb{P})}\left(\sqrt{V(\rho;H_{S})}+\sqrt{V(\tilde{\Lambda}(\rho);H_{S^{\prime}})}\right).$ (S.57)
Therefore, we only have to evaluate $\delta(\tilde{\Lambda},\mathbb{P})$, ${\cal C}(\tilde{\Lambda},\mathbb{P})$ and $V(\tilde{\Lambda}(\rho);H_{S^{\prime}})$.
In the same manner as the proof of Theorem S.2, we obtain
$\delta(\tilde{\Lambda},\mathbb{P})\leq\epsilon.$ (S.58)
Also, due to (S.49), the following is valid for arbitrary $\rho\in\mathrm{span}(\mathbb{P})$:
$\sqrt{V(\tilde{\Lambda}(\rho);H_{S^{\prime}})}$ $\leq\sqrt{V(\Lambda(\rho);H_{S^{\prime}})+\epsilon\Delta^{2}}\leq\sqrt{V(\Lambda(\rho);H_{S^{\prime}})}+\sqrt{\epsilon}\Delta.$ (S.59)
Therefore, to obtain (S.50), we only have to show
${\cal C}(\tilde{\Lambda},\mathbb{P})\geq{\cal C}(\Lambda,\mathbb{P})-2\epsilon\Delta.$ (S.60)
In the same manner as (S.38), we obtain for =$\mathbb{P}=\{|{\psi}\rangle\!\langle{\psi}|,|{\phi}\rangle\!\langle{\phi}|\}$ that
${\cal C}(\tilde{\Lambda},\mathbb{P})$ $\geq{\cal C}(\Lambda,\mathbb{P})-\|\sqrt{|{\psi}\rangle\!\langle{\psi}|}(\Lambda^{\dagger}(H_{S^{\prime}})-\tilde{\Lambda}^{\dagger}(H_{S^{\prime}}))\sqrt{|{\phi}\rangle\!\langle{\phi}|}\|_{2}$ (S.61)
$={\cal C}(\Lambda,\mathbb{P})-|\langle\psi|\tilde{\Lambda}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle|.$
The second term can be upper bounded as
$|\langle\psi|\tilde{\Lambda}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle|$ $\leq\frac{1}{2}|\langle\psi|\tilde{\Lambda}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle+\langle\phi|\tilde{\Lambda}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\psi\rangle|$ (S.62)
$\quad+\frac{1}{2}|\langle\psi|\tilde{\Lambda}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle-\langle\phi|\tilde{\Lambda}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\psi\rangle|$
$=\frac{1}{2}\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(|{\psi}\rangle\!\langle{\phi}|+|{\phi}\rangle\!\langle{\psi}|))\right|+\frac{1}{2}\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(|{\psi}\rangle\!\langle{\phi}|-|{\phi}\rangle\!\langle{\psi}|))\right|$
$=\frac{1}{2}\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(\eta_{+}-\eta_{-}))\right|+\frac{1}{2}\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(\eta^{\prime}_{+}-\eta^{\prime}_{-}))\right|$
$\leq\frac{1}{2}\left(\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(\eta_{+}))\right|+\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(\eta_{-}))\right|\right.$
$\enskip+\left.\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(\eta^{\prime}_{+}))\right|+\left|\mathrm{Tr}(H_{S^{\prime}}(\tilde{\Lambda}-\Lambda)(\eta^{\prime}_{-}))\right|\right)$
$\leq 2\epsilon\Delta$
where we introduced $\eta_{\pm}=|{\eta_{\pm}}\rangle\!\langle{\eta_{\pm}}|$ and $\eta_{\pm}^{\prime}=|{\eta_{\pm}^{\prime}}\rangle\!\langle{\eta_{\pm}^{\prime}}|$ with $|\eta_{\pm}\rangle\coloneqq\frac{1}{\sqrt{2}}(|\psi\rangle\pm|\phi\rangle)$ and $|\eta_{\pm}^{\prime}\rangle\coloneqq\frac{1}{\sqrt{2}}(|\psi\rangle\pm i|\phi\rangle)$. Therefore, we have obtained (S.60), showing (S.50). ∎
@@ II.3 Towards the extension to systems with unbounded Hamiltonians
III Construction of pairwise reversible Gibbs-preserving Operations (Proof of Theorem 2)
@@
$\langle i|\psi|i\rangle\neq 0,\quad\langle j|\psi|j\rangle\neq 0,\quad E_{i,S}\neq E_{j,S}$ (S.65)
@@
Let $\psi$ and $\phi$ be the pure states in (S.65) and (S.66). Define
$\Lambda(\rho)=\mathrm{Tr}(\psi\rho)\phi+\mathrm{Tr}\left[(\mathds{1}-\psi)\rho\right]\eta$ (S.67)
@@
$\eta=\frac{\tau_{S^{\prime}}-\mathrm{Tr}(\psi\tau_{S})\phi}{\mathrm{Tr}\left[(\mathds{1}-\psi)\tau_{S}\right]}.$ (S.68)
@@
$\phi\leq 2^{D_{\max}(\phi\|\tau_{S^{\prime}})}\tau_{S^{\prime}}=2^{D_{\min}(\psi\|\tau_{S})}\tau_{S^{\prime}}=\mathrm{Tr}(\psi\tau_{S})^{-1}\tau_{S^{\prime}}$ (S.69)
where the first inequality is by definition of $D_{\max}$, the first equality is due to (S.66), and the last equality is by definition of $D_{\min}$. This ensures that $\eta\geq 0$ and that $\Lambda$ is a valid Gibbs-preserving channel.
@@
$\mathrm{Tr}(\phi\eta)\propto\mathrm{Tr}(\phi\tau_{S^{\prime}})-\mathrm{Tr}(\psi\tau_{S})=2^{-D_{\min}(\phi\|\tau_{S^{\prime}})}-2^{-D_{\min}(\psi\|\tau_{S})}=0$ (S.70)
where the final equality is due to (S.66). This implies that $\phi$ and $\eta$ are perfectly distinguishable and thus ensures that $\delta(\Lambda,\mathbb{P})=0$ for a choice of state pair =$\mathbb{P}=\{\psi,\sigma\}$ where $\sigma$ is an arbitrary state in $S$ such that $\mathrm{Tr}(\psi\sigma)=0$.
As a choice of $\sigma$, we can particularly choose an orthogonal pure state $\psi^{\perp}$ defined as follows. Let $a_{i}\coloneqq\langle{i}|{\psi}\rangle$ and $a_{j}\coloneqq\langle{j}|{\psi}\rangle$ be the coefficients for $i$ th and $j$ th energy of $\psi$, where $i$ and $j$ are the labels in (S.65). We then choose
$|\psi^{\perp}\rangle=\frac{1}{\sqrt{|a_{j}|^{2}+|a_{i}|^{2}}}\left(a^{*}_{j}|i\rangle-a^{*}_{i}|j\rangle\right),$ (S.71)
for which one can directly check that the condition $\mathrm{Tr}(\psi\psi^{\perp})=0$ is satisfied. Here, $a^{*}_{i,j}$ is the complex conjugate of the coefficients $a_{i,j}$. Since
$\Lambda^{\dagger}(H_{S^{\prime}})=\mathrm{Tr}(\phi H_{S^{\prime}})\psi+\mathrm{Tr}(\eta H_{S^{\prime}})(\mathds{1}-\psi),$ (S.72)
@@
$\langle{\psi}|{\Lambda^{\dagger}(H_{S^{\prime}})}|{\psi^{\perp}}\rangle=0.$ (S.73)
@@
$\left|\langle{\psi}|{H-\Lambda^{\dagger}(H_{S^{\prime}})}|{\psi^{\perp}}\rangle\right|$ $=\left|\langle{\psi}|{H}|{\psi^{\perp}}\rangle\right|={\color[rgb]{0,0,0}\frac{|a^{*}_{i}a^{*}_{j}|}{\sqrt{|a_{i}|^{2}+|a_{j}|^{2}}}|E_{i,S}-E_{j,S}|}>0$ (S.74)
@@
${\cal C}(\Lambda,\mathbb{P})=\left|\langle{\psi}|{H-\Lambda^{\dagger}(H_{S^{\prime}})}|{\psi^{\perp}}\rangle\right|>0$ (S.75)
@@
$\tau_{S^{\prime},i^{\prime}}=r\tau_{S,i}+(1-r)\tau_{S,j},$ (S.77)
whose existence is guaranteed because of (S.76). Let $\psi$ and $\phi$ be the pure states defined by
$|\psi\rangle$ $:=\sqrt{r}|i\rangle_{S}+\sqrt{1-r}|j\rangle_{S},$ (S.78)
$|\phi\rangle$ $:=|i^{\prime}\rangle_{S^{\prime}}.$ (S.79)
It suffices to show that $\psi$ and $\phi$ satisfy (S.65) and (S.66). It is straightforward to check (S.65) noting $10<r<$ 1. Eq. (S.66) can be checked as follows:
@@
$\Lambda(\rho)=\mathrm{Tr}(|{+}\rangle\!\langle{+}|_{i,i+1}\rho)|{0}\rangle\!\langle{0}|+\mathrm{Tr}(|{-}\rangle\!\langle{-}|_{i,i+1}\rho)|{1}\rangle\!\langle{1}|+\mathrm{Tr}(P_{i,i+1}^{\perp}\rho)\eta$ (S.81)
@@
$\eta\coloneqq\frac{\tau_{S^{\prime}}-\mathrm{Tr}(|{+}\rangle\!\langle{+}|_{i,i+1}\tau_{S})|{0}\rangle\!\langle{0}|-\mathrm{Tr}(|{-}\rangle\!\langle{-}|_{i,i+1}\tau_{S})|{1}\rangle\!\langle{1}|}{\mathrm{Tr}(P_{i,i+1}^{\perp}\tau_{S})}.$ (S.82)
@@
$\eta\propto\left(\tau_{S^{\prime},0}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\right)|{0}\rangle\!\langle{0}|+\left(\tau_{S^{\prime},1}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\right)|{1}\rangle\!\langle{1}|+\sum_{j\geq 2}\tau_{S^{\prime},j}|{j}\rangle\!\langle{j}|\geq 0$ (S.83)
@@
$\tau_{S^{\prime},k}-\tau_{S,k}=\frac{(Z-Z^{\prime})e^{-\beta E_{k}}}{ZZ^{\prime}}\geq 0,\quad Z\coloneqq\sum_{n=0}^{d-1}e^{-\beta E_{n}},Z^{\prime}\coloneqq\sum_{n=0}^{d^{\prime}-1}e^{-\beta E_{n}}$ (S.84)
@@
$\tau_{S^{\prime},0}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\geq\tau_{S,0}-\frac{\tau_{S,i}+\tau_{S,i+1}}{2}\geq 0$ (S.85)
@@
$\Lambda^{\dagger}(H_{d^{\prime}})=\mathrm{Tr}(H_{d^{\prime}}|{0}\rangle\!\langle{0}|)|{+}\rangle\!\langle{+}|_{i,i+1}+\mathrm{Tr}(H_{d^{\prime}}|{1}\rangle\!\langle{1}|)|{-}\rangle\!\langle{-}|_{i,i+1}+\mathrm{Tr}(H_{d^{\prime}}\eta)P_{i,i+1}^{\perp}$ (S.86)
@@
+|${}_{i,i+1}\langle{+}|{\Lambda^{\dagger}(H_{d^{\prime}})}|{-}\rangle_{i,i+1}=0.$ (S.87)
@@
|+$\left|{}_{i,i+1}\langle{+}|{H_{d}}|{-}\rangle_{i,i+1}\right|=\frac{E_{i+1}-E_{i}}{2}>0$ (S.88)